# On the Hyperparameter Loss Landscapes of Machine Learning Algorithms

### v1

@@ B.2 Landscape Analysis Metrics
Number of Local Optima. A configuration $\bm{\lambda}$ is said to be a local optimum if its performance is superior to any other configuration in its neighborhood, i.e., $\forall\bm{\lambda}\in{\mathcal{N}(\bm{\lambda}^{\ell})}$, we have $\mathcal{L}(\bm{\lambda}^{\ell})<{\mathcal{L}(\bm{\lambda})}$. For a unimodal landscape, there is only a global optimum configuration $\bm{\lambda}^{*}$. In constrast, multimodal landscapes have various local optima with sub-optimal performance, which can pose challenges to the optimization.
@@
B.3 Local Optima Network
Beyond the number of local optima and their basin sizes, a further aspect to investigate is the connectivity pattern between them. For example, an important question that we may concern is whether we can escape from a given local optimum to the global optimum?, if yes, then what is the chance of this? Local optima networks (LON) (Ochoa et al., 2008; Vérel et al., 2011), which are rooted in the study of energy landscapes in chemical physics (Stillinger, 1995), address these questions by constructing a subspace of the original fitness landscape where the nodes indicate local optima, and edges represent possible transitions between them. In particular, an improving edge can be traced from local optimum configuration $\bm{\lambda}^{\ell}_{i}$ to $\bm{\lambda}^{\ell}_{j}$ if configurations in $\mathcal{B}(\bm{\lambda}^{\ell}_{i})$ can escape to $\bm{\lambda}^{\ell}_{j}$ by applying a $2$-bit perturbation followed by local search. The edge weights $w_{i,j}$ indicate the total probability for such transitions to happen (see Algorithm 3) between two local optima. By conducting network mining on LONs (e.g., Huang & Li (2023)), we can get further insights into the distribution and connection between local optima as well as how they are potentially linked with the global optimum. However, while we believe LON could be an effective tool for analyzing complex landscapes and we incorporate this as part of our landscape analysis tool, we do not present such analyses in most scenarios in this paper since their HP loss landscapes only possess a few local optima (if not none).
@@ B.3 Local Optima Network
0:  The set of local optima $\mathcal{V}$; The basin of attraction of each local optimum $\mathcal{B}_{\mathbf{c}^{\ell}}$; The set of all configurations $\mathcal{C}$; A neighborhood function $\mathcal{N}_{d}(\mathbf{c})$
1:  $\mathcal{E}\leftarrow\emptyset$
2:  $\mathcal{W}\leftarrow\emptyset$
3:  for all $\mathbf{c}^{\ell}\in\mathcal{V}$ do
4:     for all $\mathbf{c}^{\ell\prime}\in\mathcal{N}_{2}(\mathbf{c}^{\ell})$ do
5:        $\mathbf{c}^{\ell}_{\text{new}}\leftarrow\textsc{LocalSearch}(\mathbf{c}^{\ell\prime})$
6:        if $f(\mathbf{c}^{\ell}_{\text{new}})<f(\mathbf{c}^{\ell})$ then
7:           if edge $(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})$ not in $\mathcal{E}$ then
8:              $\mathcal{E}\leftarrow\mathcal{E}\cup\{(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})\}$
9:              $\mathcal{W}[(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})]\leftarrow 1$
10:           else
11:              $\mathcal{W}[(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})]\leftarrow\mathcal{W}[(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})]+1$
12:           end if
13:        end if
14:     end for
15:  end for
16:  return $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{W})$
@@ B.4 Landscape Similarity Metrics
Shake-up Metric. It originates from the Kaggle competition community and is designed to assess the rank changes between the public and private leaderboards of a competition (Trotman, 2019). To be specific, it quantifies the average movement of rankings from public board to the private board. For HP loss landscapes, this metric can indicate the expected rank shifts for a configuration when evaluating it on two different scenarios (e.g., change the dataset).
@@
$\gamma$-set Similarity. It is proposed in Watanabe et al. (2023a) to assess the similarity of two tasks using the ratio of their most proninent configurations. More specifically, for two HP loss landscapes, their $\gamma$-set similarity is defined as the ratio of the intersection of the top-$\gamma$ regions to the union of them. In this paper, we consider $\gamma=10\%$.
Appendix C Analysis on NAS Loss Landscapes
C.1 NAS Benchmarks
@@
NAS-Bench-101. It represents the first effort towards benchmarking NAS research and thus fostering reproducibility in the community. It evaluates a CNN architecture with $3$ stacked blocks, where a down-sampling is added between two consecutive blocks; A $3\times 3$ convolution is used before the main blocks, and the outputs of the main blocks are fed to an average pooling and fully connected layer. NAS-Bench-101 considers a cell-based search space which enumerates all possible configurations for each block. More specifically, the search space is formulated as a directed acylic graph (DAG) with $7$ nodes and a maximum of $9$ edges. Here, each node can represent one of the following operations: $a)$: $1\times 1$ convolution, $b)$: $3\times 3$ convolution, and $c)$: max pooling. After removing all isomorphic cells, this search space results in $423$k unique configurations. NAS-Bench-101 evaluates each of them on the CIFAR-10 dataset and records meta-data at the $\{4,12,36,108\}^{\text{th}}$ epoch.
@@
NAS-Bench-201. It features a different skeleton compared to NAS-Bench-101, in which a residual block is applied to connect $3$ cells. Each cell here is a DAG with $4$ nodes and $6$ edges. Morever, here, operations are represented by edges, which have $5$ types: $a)$: zeroize (do nothing), $b)$: $1\times 1$ convolution, $c)$: $3\times 3$ convolution, $d)$: $3\times 3$ average pooling, $e)$: skip connection. The benchmark thus contains $5^{6}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}15,625}$ unique model architectures, with each evaluated on $3$ different datasets: $i)$: CIFAR-10, $ii)$: CIFAR-100, $iii)$: ImageNet-16-120 using $200$ epochs.
C.2 NAS Landscape Construction
@@
NAS-Bench-101 Neighborhood. In the original paper, a cell-encoding method based on adjacency matrices is introduced to encode configurations, which comprises two components. First, a $7\times 7$ upper-triangular binary matrix is used to indicate whether an edge exist between two nodes and thus determine the connectivity pattern of nodes. Next, the functionality of a configuration also depends on which operation is performed at each node, and this could be encoded via a vector of length $5$ (the input and output nodes are the same across archiectures and are omitted, non-existent nodes are represented by NaN). Therefore, a configuration could be specified using an adjacency matrice and a node vector. We define a neighbor of a given configuration to be the one with $1$-edit distance from it (e.g., either adding or deleting one edge, or change the operation of one node), as in the original paper. Note that not all such configurations are valid in the benchmark, as they can be isomorphic to others. The benchmark API provides built-in function to check for this.
@@
NAS-Bench-201 Neighborhood. The encoding of configurations in this benchmark is much more straight forward. Specifically, we encode each configuration using a $6$-bit vector, where each bit specifies the operation taken at each edge, which can take $5$ categorical values. We base our neighborhood definition here on the $1$-edit distance as well, in which a $1$-bit mutant of a configuration is the one with the operation in only one edge altered.
C.3 Results on NAS Loss Landscapes
@@
While we have conducted all analyses on both benchmarks, it would be too tedious to lay all the information here. Instead, we use NAS-Bench-101 to discuss general NAS landscape characteristics and multi-fidelity, and present the comparisons between NAS landscapes across datasets using NAS-Bench-201. Before we present our results, we also note that while the majority of configurations ($359$k out of $423$k) in NAS-Bench-101 search space come with $7$ nodes (i.e., with $5$ intermediate operations), the rest of them have less nodes. Here we mainly focus on configuragtions with $7$ nodes, since accounting all the configurations would result in many independent components in the landscape. We do not expect this to significantly affect our results.
@@
NAS Landscape Visualization. We first visulize the NAS-Bench-101 landscape using our proposed landscape visualization method in Section 2, as shown in Figure 8. Specifically, we plot the training accuracy landscape along with test accuracy landscapes trained at $4$ different number of epochs. It is clear to see from the plot that the landscape is far from unimodal, with many local optima. However, configurations still tend to form local clusters, though the relative size of each plateau seems to be much smaller than we see for HP loss landscapes. Considering that the search space here contains nearly $30$ times more confiurations than the HP space we used in the main text, each cluster may contain thousands of configurations, inside which the landscape could still be sufficiently smooth.
Figure 8: Visualization of NAS-Bench-101 landscape using our proposed method. Here, we present the training accuracy landscapes as well as test accuracy landscapes at different recorded epochs. Color indicates rank of test accuracy, and higher values are better.
@@
NAS Landscape Metrics. Here we report several landscape metrics for the NAS-Bench-101 landscape with respect to test accuracy at the $108$-th epoch:
@@
  • Number of Local Optima. One of the most distinguishable property of NAS-Bench-101 is a large collection of local optima in the landscape. In fact, we found $5,908$ local optima in total (out of the $359$k configurations), which could probably make the landscape far more difficult to optimization. We will discuss more about them and their basins in the LON part.
Figure 9: Visualization of the local optima network of NAS-Bench-101 test landscapes at the $108$-th epoch, containing $5,908$ nodes. Radius of each node (i.e., local optimum) indicates size of the corresponding basin of attraction. The color indicates test accuracy, where warmer color is better. Edges indicate transition probabilities between local optima, where thicker, warmer edges imply the corresponding transition if more likely to happen. Edge directions indicate the improving direction (i.e., pointing to the fitter configuration).
NAS Landscape Across Fidelities. From Figure 8, we could see that in general the test landscapes with lower fidelities resemble the one trained with $108$ epoch. Quantitatively, the Spearman correlation between the test accuracy landscapes at $108$-th epoch and $32$-th epoch is $0.904$, with a Shake-up metric of $9.74\%$. These suggest a good general correlation between these landscapes, although there are $4$ times difference in their budget. When zooming into the $10\%$ region, the $\gamma$-set similarity is $0.64$, which is also farily good (the intesection ratio between top-$10\%$ regions is $78\%$). When we further decrease the budge by $4$ times, the Spearman correlation and Shake-up metric obtained between the $12$-th and $108$-th epoch are $0.657$ and $18.4\%$ respectively, while the $\gamma$-set similarity is $0.317$. Finally, with only $4$ epochs of training, the above metrics further change to $0.504$, $22.4\%$ and $0.164$ respectively. In general, the landscapes are still moderately correlated, but the detailed pattern can be largely distorted.
Figure 10: Visualization of NAS-Bench-201 landscape using our proposed method across datasets.
Local Optima Network Analysis.
@@
Figure 9 shows the local optima network of NAS-Bench-101 test landscapes at the $108$-th epoch. It could be obviously seen that there are clear community (clustering) structure in the network, where local optima with large basin of attractions tend to locate at the center of each cluster, which usually feature a promising performance. To be specific, from the left plots in Figure 11, we could clearly observe that the size and radius [^5] Here, we define the radius of a basin of attraction as the expected number of local search steps needed to reach the corresponding local optimum. Since in practice, we observe that this value is often highly correlated with the size of basin of attraction, and correlate with performance in a similar manner, we only report basin size in the main text. of the basin of attraction is positively correlated with the performance of local optima (Spearman correlation $>0.55$). More importantly, as suggested by the cumulative distribution of basin size shown in Figure 12, the highly-fit local optima have a dominant size of basin of attractions in the landscape. For example, the dashed line in Figure 12 indicates that the cumulative sum of basin sizes of those local optima with $acc>94.3\%$ take $50\%$ of the total basin size (which is the number of total configurations in the landscape). Since being in the basin of a local optimum would imply that local search will eventually converge it, our result is to say that if we start local search from a random configuration in the landscape, there is $50\%$ chance that we would end up in a local optimum with $acc>94.3\%$. This is a very promising result, since we find such a level of accuracy is already better than $98.58\%$ of total configurations in the whole search space! It is also superior than $76.84\%$ of other local optima in the landscape.
@@
Beyond local search, we then continue to consider if we do fall in a local optima whose fitness is not that satisfactory, then what is the chance that we can somehow escape from it?. The two plots at the right panel of Figure 11 show the distribution of escape rate and improve rate correlate with test accuracy. Here, the escape rate is the chance that, a configuration in the basin of a local optimum $\mathbf{c}^{\ell}$, after applying a $2$-bit perturbation, would converge to a different local optimum $\mathbf{c}^{\ell}_{\text{new}}$. On the other hand, the improve rate further restricts that the new local optimum should have better performance compared to the current one. From Figure 11, we could clearly see observe the majority of the local optima feature a escape rate of larger than $50\%$, which suggests that most local optima are not that difficult to escape from. For improve rate, we observe a good correlation with test accuracy, where it is more easy to find an improving move for a poorly-performed local optimum. Unfortunately, for those that already have promising performance, there is only little chance to transit to a better basin using $2$-bit perturbation.
C.4 Results on NAS-Bench-201
We visualize the loss landscapes of NAS-Bench-201 on the $3$ datasets, namely CIFAR-10/100 and ImageNet, in Figure 10. In general, we see that results on these $3$ tasks reveal strong consistency to each other (Spearman correlation $>0.95$), which conforms with our findings on HP loss landscapes.
Appendix D Details About FCNet HP Loss Landscapes
Figure 13: Local optima network for the $100$ local optima in the FCNet landscape on parkinsons telemonitoring dataset. The left plot shows the full-view of the network, where node size represents the size of the corresponding basin of attraction. Node color indicates the performance of the model, which is labeled as rank values for each node, and the global optimum has a rank of $1$. The right plot shows the neighborhoods of the global optimum, while other nodes are omitted.
Table 3: Landscape Features of FCNet on Different Datasets
 Dataset Autocorr. Assortativity Neutrality NDC #LO Mean Basin Size
Parkinsons Tele. $0.4683$ $0.6232$ $0.4317$ $0.6368$ $100$ $589$
Protein Struc. $0.4896$ $0.5789$ $0.4746$ $0.7411$ $128$ $385$
Naval Prop. $0.5720$ $0.5298$ $0.4091$ $0.6433$ $347$ $159$
Slice Local. $0.4887$ $0.6134$ $0.4427$ $0.8216$ $24$ $2372$
D.1 Landscape Features
Figure 14: Spearman correlation between LON features ($i)$: mean basin size, $ii)$: mean basin radius, $iii)$: maximum basin radius, $iv)$: in-coming strengths, $v)$: out-going strengths, $vi)$: closeness centrality, $vii)$: eigenvector centrality, $viii)$: pagerank centrality) and local optima performance for FCNet across datasets.
From the visualization of FCNet landscapes shown in Figure 2, we can conclude that they generally follow the same characteristics as we discussed in Section 4.1. Most of such observations are supported by FLA metrics reported in Table 3. However, we note that all four landscapes seem to reveal multimodal structures, as there are dozens to hundreds of local optima with relative large basins (up to a mean size of $2,372$). Similar observations have been made by Pushak & Hoos, who speculated that the reasons could be FCNet scenarios fall into the over-parameterized regime, which is different from the other scenarios. While we also agree with this reasoning, we seek to conduct further analysis on the local optima using the local optima network (LON) (Appendix B.3). Figure 13 (left) shows the LON of FCNet on the parkinsons telemonitoring dataset. It can be seen that in general, local optima with better performance tend to feature a large basin of attraction and high in-coming strenghts (i.e., in LON context, the total probability for other local optima to reach it). This implies that, despite the pressence of many other local optima, the global optima still plays a pivotal role in the connectivity of the network (see Figure 13, right), which is not the worst case that one may encounter (e.g., a global optimum with tiny basin locates at a secluded region that the optimizer can hardly find.)
Appendix E Landscape Similarity & Datasets Characteristics

### v2

@@ B.2 Landscape Analysis Metrics
Number of local optima. A configuration $\bm{\lambda}$ is said to be a local optimum if its performance is superior to any other configuration in its neighborhood, i.e., $\forall\bm{\lambda}\in{\mathcal{N}(\bm{\lambda}^{\ell})}$, we have $\mathcal{L}(\bm{\lambda}^{\ell})<{\mathcal{L}(\bm{\lambda})}$. For a unimodal landscape, there is only a global optimum configuration $\bm{\lambda}^{*}$. In constrast, multimodal landscapes have various local optima with sub-optimal performance, and is known to be ‘rugged’.
@@
B.3 Local Optima Network (LON)
Beyond the number of local optima and their basin sizes, a further landscape property to interrogate is the distribution of local optima and the connectivity pattern between them. For example, an important question that we may concern is whether we can escape from a given local optimum to the global optimum?, if yes, then, what is the chance of this? Local optima networks (LON) [29, 85], which are rooted in the study of energy landscapes in chemical physics [86], address these questions by constructing a subspace of the original fitness landscape where the nodes indicate local optima, and edges represent possible transitions between them. In particular, an improving edge can be traced from local optimum configuration $\bm{\lambda}^{\ell}_{i}$ to $\bm{\lambda}^{\ell}_{j}$ if configurations in $\mathcal{B}(\bm{\lambda}^{\ell}_{i})$ can escape to $\bm{\lambda}^{\ell}_{j}$ by applying a $2$-bit perturbation followed by local search. The edge weights $w_{i,j}$ indicate the total probability for such transitions to happen (see Algorithm 3) between two local optima. By conducting network mining on LONs (e.g., [21]), we can get further insights into the distribution and connection between local optima as well as how they are potentially linked with the global optimum.
0:  The set of local optima $\mathcal{V}$; The basin of attraction of each local optimum $\mathcal{B}_{\mathbf{c}^{\ell}}$; The set of all configurations $\mathcal{C}$; A neighborhood function $\mathcal{N}_{d}(\mathbf{c})$ 1:  $\mathcal{E}\leftarrow\emptyset$ 2:  $\mathcal{W}\leftarrow\emptyset$ 3:  for all $\mathbf{c}^{\ell}\in\mathcal{V}$ do 4:    for all $\mathbf{c}^{\ell\prime}\in\mathcal{N}_{2}(\mathbf{c}^{\ell})$ do 5:       $\mathbf{c}^{\ell}_{\text{new}}\leftarrow\textsc{LocalSearch}(\mathbf{c}^{\ell\prime})$ 6:       if $f(\mathbf{c}^{\ell}_{\text{new}})<f(\mathbf{c}^{\ell})$ then 7:          if edge $(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})$ not in $\mathcal{E}$ then 8:             $\mathcal{E}\leftarrow\mathcal{E}\cup\{(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})\}$ 9:             $\mathcal{W}[(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})]\leftarrow 1$ 10:          else 11:             $\mathcal{W}[(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})]\leftarrow\mathcal{W}[(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})]+1$ 12:          end if 13:       end if 14:    end for 15:  end for 16:  return $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{W})$
@@ B.4 Landscape Similarity Metrics
Shake-up metric. It originates from the Kaggle competition community and is designed to assess the rank changes between the public and private leaderboards of a competition [63]. To be specific, it quantifies the average movement of rankings from public board to the private board. For HP loss landscapes, this metric can indicate the expected rank shifts for a configuration when evaluating it on two different scenarios (e.g., change the dataset).
@@
$\gamma$-set similarity. It is proposed in [64] to assess the similarity of two tasks using the ratio of their most proninent configurations. More specifically, for two HP loss landscapes, their $\gamma$-set similarity is defined as the ratio of the intersection of the top-$\gamma$ regions to the union of them. In this paper, we consider $\gamma=10\%$.
Appendix C HP Loss Landscapes of SVM
Figure 8: HP loss landscapes of SVM on four tabular datasets, with $\alpha=\{0.1,1\}$ training data, with respect to training and test accuracy.
Due to the inherent scalability issue of SVM, it is infeasible to evaluate it on most of our considered tabular datasets, as they often contains a large sample size and a high-dimensional feature space. Instead, we only conducted experiments on a few datasets that are relatively small in size with respective to two HPs (gamma and c) using RBF kernel. Since the resulting HP loss landscapes are already in a 2D space, we thus directly visualize them as heatmaps. Fig. 8 presents the results on four tabular datasets (#44122, #44123, #44089, #44125) with $\alpha=\{0.1,1\}$ training data. We can observe that the patterns here generally match our previous findings on other models, i.e., the landscapes are smooth, highly clustered, and nearly unimodal; a considerable portion of the landscape is neutral, and the performance difference diminishes around the optimum. These general characteristics also largely preserves across datasets, or with lower training fidelities. Nevertheless, similar to our observed patterns for XGBoost, here the training loss landscape exhibits very different topography compared to the test loss landscape, where overfitting is prevalent.
Appendix D Additional Results on Landscape Visualization.
D.1 Landscape Visualizations for DT, LGBM and RF
We present additional landscape visualizations for decision tree (DT), LightGBM (LGBM), and random forest (RF) in Fig. 9.
Figure 9: 2D visualization of HP loss landscapes of DT, LightGBM and RF under different scenarios: (A) $\mathcal{L}_{\text{test}}$ landscape on baseline datasets (#44059), (B) $\mathcal{L}_{\text{train}}$ landscape on baseline datasets, (C) Low-fidelity $\mathcal{L}_{\text{test}}$ landscape on baseline datasets obtained using $10\%$ training data, (D) $\mathcal{L}_{\text{test}}$ landscape on different datasets (#44143). Colors indicate ranks of performance (lower rank values are better).
D.2 Landscape Visualizations for NK-Landscape and BBOB Functions
Figure 10: Visualizations of Kauffman’s $NK$-landscape of $n=10$ with different levels of ruggedness using our proposed method in Section 2 (low: $k=0$, medium: $k=5$, and strong: $k=10$).
Figure 11: Recovered landscapes with our proposed visualization method versus groundtruth landscape (reference) surfaces of $f_{1}$ and $f_{17}$ to $f_{24}$ in the black-box optimization benchmark (BBOB) benchmarking suite from COCO.
In order to demonstrate the capability of our proposed landscape visualization method in capturing the structural characteristics of fitness landscapes, here we conduct additional experiments with artifical black-box optimization problems.
NK-landscape. We first consider the well-known $NK$-landscape model [74], which was introduced by Stuart Kauffman in the last century for studying the evolutionary of biological systems. It has two tunable parameters: $n$ and $k$. Here, $n$ represents the number of elements in the system, while $k$ determines the level of interaction between these elements. As $k$ increases, the landscape becomes increasingly rugged, with a greater number of local optima. For illustration here, we generate $NK$-landscapes with $n=10$ and three levels of ruggedness by varying the parameter $k\in\{0,5,10\}$. As shown in Fig. 10, our visualization method is able to capture the increasing ruggedness of the landscapes as $k$ grows, where the landscape becomes more rugged.
BBOB benchmark suite. Another scenario we consider here is the black-box optimization benchmark (BBOB) functions suite from the COCO platform [75]. This suite contains $24$ BBOB functions ($f_{1}$ to $f_{24}$) with a diverse range of complexities that are widely studied in literature. Here we test the capability of our visualization method in recovering the topography of these functions by comparing the visualized landscapes with the groundtruth surfaces. In particular, we present the results for $f_{17}$ to $f_{24}$ in Fig. 11, which are known to exhibit strong multimodality (to compare, we also present $f_{1}$, a highly separable unimodal function). As can be seen from the results, our method is able to accurately recover shape of the function surface, even for the most complicated ones.
Appendix E Analysis on NAS Loss Landscapes
E.1 NAS Benchmarks
@@
NASBench-101. It represents the first effort towards benchmarking NAS research and thus fostering reproducibility in the community. It evaluates a CNN architecture with $3$ stacked blocks, where a down-sampling is added between two consecutive blocks; A $3\times 3$ convolution is used before the main blocks, and the outputs of the main blocks are fed to an average pooling and fully connected layer. NASBench-101 considers a cell-based search space which enumerates all possible configurations for each block. More specifically, the search space is formulated as a directed acylic graph (DAG) with $7$ nodes and a maximum of $9$ edges. Here, each node can represent one of the following operations: $a)$: $1\times 1$ convolution, $b)$: $3\times 3$ convolution, and $c)$: max pooling. After removing all isomorphic cells, this search space results in $423$k unique configurations. NASBench-101 evaluates each of them on the CIFAR-10 dataset and records meta-data at the $\{4,12,36,108\}^{\text{th}}$ epoch.
@@
NASBench-201. It features a different skeleton compared to NASBench-101, in which a residual block is applied to connect $3$ cells. Each cell here is a DAG with $4$ nodes and $6$ edges. Morever, here, operations are represented by edges, which have $5$ types: $a)$: zeroize (do nothing), $b)$: $1\times 1$ convolution, $c)$: $3\times 3$ convolution, $d)$: $3\times 3$ average pooling, $e)$: skip connection. The benchmark thus contains $5^{6}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}15,625}$ unique model architectures, with each evaluated on $3$ different datasets: $i)$: CIFAR-10, $ii)$: CIFAR-100, $iii)$: ImageNet-16-120 using $200$ epochs.
[…]
E.2 NAS Landscape Construction
@@
NASBench-101 Neighborhood. In the original paper, a cell-encoding method based on adjacency matrices is introduced to encode configurations, which comprises two components. First, a $7\times 7$ upper-triangular binary matrix is used to indicate whether an edge exist between two nodes and thus determine the connectivity pattern of nodes. Next, the functionality of a configuration also depends on which operation is performed at each node, and this could be encoded via a vector of length $5$ (the input and output nodes are the same across archiectures and are omitted, non-existent nodes are represented by NaN). Therefore, a configuration could be specified using an adjacency matrice and a node vector. We define a neighbor of a given configuration to be the one with $1$-edit distance from it (e.g., either adding or deleting one edge, or change the operation of one node), as in the original paper. Note that not all such configurations are valid in the benchmark, as they can be isomorphic to others. The benchmark API provides built-in function to check for this.
@@
NASBench-201/360 Neighborhood. The encoding of configurations in this search space is much more straight forward. Specifically, we encode each configuration using a $6$-bit vector, where each bit specifies the operation taken at each edge, which can take $5$ categorical values. We base our neighborhood definition here on the $1$-edit distance as well, in which a $1$-bit mutant of a configuration is the one with the operation in only one edge altered.
E.3 Results on NAS Loss Landscapes
@@
While we have conducted all analyses on both benchmarks, it would be too tedious to lay all the information here. Instead, we use NASBench-101 to discuss general NAS landscape characteristics and multi-fidelity, and present the comparisons between NAS landscapes across datasets using NASBench-201. Before we present our results, we also note that while the majority of configurations ($359$k out of $423$k) in NASBench-101 search space come with $7$ nodes (i.e., with $5$ intermediate operations), the rest of them have less nodes. Here we mainly focus on configuragtions with $7$ nodes, since accounting all the configurations would result in many independent components in the landscape. We do not expect this to significantly affect our results.
@@
NAS Landscape Visualization. We first visulize the NASBench-101 landscape using our proposed landscape visualization method in Section 2, as shown in Fig. 2 (C). Specifically, we plot the training accuracy landscape along with test accuracy landscapes trained at $4$ different number of epochs. It is clear to see from the plot that the landscape is far from unimodal, with many local optima. However, configurations still tend to form local clusters, though the relative size of each plateau seems to be much smaller than we see for HP loss landscapes. Considering that the search space here contains nearly $30$ times more confiurations than the HP space we used in the main text, each cluster may contain thousands of configurations, inside which the landscape could still be sufficiently smooth.
@@
NAS Landscape Metrics. Here we report several landscape metrics for the NASBench-101 landscape with respect to test accuracy at the $108$-th epoch:
@@
  • Number of Local Optima. One of the most distinguishable property of NASBench-101 is a large collection of local optima in the landscape. In fact, we found $5,908$ local optima in total (out of the $359$k configurations), which could probably make the landscape far more difficult to optimization. We will discuss more about them and their basins in the LON part.
Figure 12: Visualization of the local optima network of NASBench-101 test landscapes at the $108$-th epoch, containing $5,908$ nodes. Radius of each node (i.e., local optimum) indicates size of the corresponding basin of attraction. The color indicates test accuracy, where warmer color is better. Edges indicate transition probabilities between local optima, where thicker, warmer edges imply the corresponding transition if more likely to happen. Edge directions indicate the improving direction (i.e., pointing to the fitter configuration).
NAS Landscape Across Fidelities. From Fig. 2 (C), we could see that in general the test landscapes with lower fidelities resemble the one trained with $108$ epoch. Quantitatively, the Spearman correlation between the test accuracy landscapes at $108$-th epoch and $32$-th epoch is $0.904$, with a Shake-up metric of $9.74\%$. These suggest a good general correlation between these landscapes, although there are $4$ times difference in their budget. When zooming into the $10\%$ region, the $\gamma$-set similarity is $0.64$, which is also farily good (the intesection ratio between top-$10\%$ regions is $78\%$). When we further decrease the budge by $4$ times, the Spearman correlation and Shake-up metric obtained between the $12$-th and $108$-th epoch are $0.657$ and $18.4\%$ respectively, while the $\gamma$-set similarity is $0.317$. Finally, with only $4$ epochs of training, the above metrics further change to $0.504$, $22.4\%$ and $0.164$ respectively. In general, the landscapes are still moderately correlated, but the detailed pattern can be largely distorted.
@@
Local Optima Network Analysis. Fig. 12 shows the local optima network of NASBench-101 test landscapes at the $108$-th epoch. It could be obviously seen that there are clear community (clustering) structure in the network, where local optima with large basin of attractions tend to locate at the center of each cluster, which usually feature a promising performance. To be specific, from the left plots in Fig. 13, we could clearly observe that the size and radius [^5] Here, we define the radius of a basin of attraction as the expected number of local search steps needed to reach the corresponding local optimum. Since in practice, we observe that this value is often highly correlated with the size of basin of attraction, and correlate with performance in a similar manner, we only report basin size in the main text. of the basin of attraction is positively correlated with the performance of local optima (Spearman correlation $>0.55$). More importantly, as suggested by the cumulative distribution of basin size shown in Fig. 14, the highly-fit local optima have a dominant size of basin of attractions in the landscape. For example, the dashed line in Fig. 14 indicates that the cumulative sum of basin sizes of those local optima with $acc>94.3\%$ take $50\%$ of the total basin size (which is the number of total configurations in the landscape). Since being in the basin of a local optimum would imply that local search will eventually converge it, our result is to say that if we start local search from a random configuration in the landscape, there is $50\%$ chance that we would end up in a local optimum with $acc>94.3\%$. This is a very promising result, since we find such a level of accuracy is already better than $98.58\%$ of total configurations in the whole search space! It is also superior than $76.84\%$ of other local optima in the landscape.
@@
Beyond local search, we then continue to consider if we do fall in a local optima whose fitness is not that satisfactory, then what is the chance that we can somehow escape from it?. The two plots at the right panel of Fig. 13 show the distribution of escape rate and improve rate correlate with test accuracy. Here, the escape rate is the chance that, a configuration in the basin of a local optimum $\mathbf{c}^{\ell}$, after applying a $2$-bit perturbation, would converge to a different local optimum $\mathbf{c}^{\ell}_{\text{new}}$. On the other hand, the improve rate further restricts that the new local optimum should have better performance compared to the current one. From Fig. 13, we could clearly see observe the majority of the local optima feature a escape rate of larger than $50\%$, which suggests that most local optima are not that difficult to escape from. For improve rate, we observe a good correlation with test accuracy, where it is more easy to find an improving move for a poorly-performed local optimum. Unfortunately, for those that already have promising performance, there is only little chance to transit to a better basin using $2$-bit perturbation.
Appendix F Landscape Similarity & Datasets Characteristics